# A Unified Sequence Parallelism Approach for Long Context Generative AI

### v1

@@ Abstract
Sequence parallelism (SP), which divides the sequence dimension of input tensors across multiple computational devices, is becoming key to unlocking the long-context capabilities of generative AI models. This paper investigates the state-of-the-art SP approaches, i.e. DeepSpeed-Ulysses and Ring-Attention, and proposes a unified SP approach, which is more robust to transformer model architectures and network hardware topology. This paper compares the communication and memory cost of SP and existing parallelism, including data/tensor/zero/expert/pipeline parallelism, and discusses the best practices for designing hybrid 4D parallelism involving SP. We achieved 86% MFU on two 8xA800 nodes using SP for sequence length 208K for the LLAMA3-8B model. Our code is publicly available on https://github.com/feifeibear/long-context-attention.
@@ 3 Unified Ulysses-Ring Sequence Parallelism
SP-Ulysses is sensitive to the number of attention heads. The parallelism degree  [^1] The parallelism degree is the number of devices that participate in parallel computation. In other words, it is the number of processes in a parallel process group. of DS-Ulysses cannot exceed the number of attention heads $hc$. Consequently, it is not suitable for GQA (Grouped Query Attention) [10] and MQA (Multi-Query Attention) [11] scenarios. For instance, Llama3-8B employs MQA with a KV head number of 8, which means that when using DS-Ulysses SP, the maximum SP degree is 8. However, if GQA is used and KV head number is 1, DS-Ulysses dose not function. In addition, since Tensor Parallelism also requires division across the $hc$ dimension, SP-Ulysses, and TP are in conflict.
@@ 4 SP in 4D Parallelism
As shown in Table 2, we analyze the communication and memory cost of a standard transformer block for different parallelism. The impact of MQA is not reflected in the table, but we will analyze it later. Entries in the Communications Params columns indicate the collective communication operations are conducted on the parameters and gradients of the transformer block. It includes the parameters/gradients of the weight and bias tensors of 4 Linear Layers in the Self-Attention layer, as well as 2 Linear Layers of the FFN layer, amounting to a volume of $12\times O(d^{2})$ elements in the GPT-2 model. Note that llama3 and llama2 has $12\times O(d^{2})$ elements since the intermediate size is lower. Entries in Communications Act columns indicate the collective communications are conducted on a single hidden states tensor belonging to activations, containing $bs\times L\times d$ elements. The memory cost is broken down into model parameters and gradients (P/G), Optimizer States (OS), and intermediate activation tensors (Act).
@@
Tip 3: We suggest that when utilizing SP, it should always be used in conjunction wit ZeRO-1. One can also consider employing ZeRO-2, ZeRO-3, and Offload techniques [5, 15] to trade off communication cost for memory savings.
@@
MQA can reduce SP communication costs, while the communication cost of TP-sp remains unchanged. Assuming the MQA group number is $G$, the Ulysses communication cost for $K$ and $V$ is reduced to 1/$G$, and the activation communication cost is reduced to $4O(bs*L*d)+4O(bs*L*d/G)$. Additionally, SP-Ring, which only communicates $K,V$, also leads to a 1/$G$ reduction in P2P bandwidth demand.
Tip 4: We suggest that SP may have an advantage over TP when employing MQA in terms of communication cost, as MQA can reduce the communication cost of SP without affecting TP.
@@ 5.1 Performance of Unified SP
As shown in Table 3, on an 8xA100-SXM4 NVLink node, the highest throughput for both 32K and 128K sequence lengths was achieved when the ulysses_degree is set to 8, that is the same as SP-Ulysses. Ulysses demonstrated a significant advantage over Ring. This verified the argument in Sec. 3 of SP-Ring, that although communication overhead can be hidden through overlapping with computation, it results in a reduction in computational efficiency.
@@ 5.2 End-to-end SP Performance in Megatron-LM
Firstly, we compare the performance of SP and DP under the same LLAMA2-7B workload on a single node of A800 GPUs. The global batch size is 8. We use the SP-Unified, and pick the best ulysses-degree and ring-degree settings. In a single node, ulysses-degree as 8 achieves the best performance. As shown in Figure 5, DP outperforms SP across various input sequence lengths, which confirms our conclusion in Tip 2.
@@ 5.4 Hybrid SP and TP
We employed a hybrid parallel strategy combining TP and SP-Unified to benchmark the throughput LLAMA3-8B training throughput in TFLOPS across two nodes. The results are shown in Table 6. The global batch size is fixed to be 1, aiming to maximize the sequence length.
Table 5 presents the performance of the llama2-7B model on a single node of 8xA800 GPU 80GB. The longest sequence length that can be achieved is 64K, and the best performance is obtained when tp-degree=2 and ulysses=2. It outperforms TP-sp-only by 10%.
When global-bs=16 and seqlen=30K, SP-Ulysses delivers the optimal performance, significantly surpassing other SP and TP hybrid strategies. It is also 26% better than TP-sp-only. This indicates that, despite the similar communication cost of PS-Ulysses and TP-sp, PS-Ulysses is more efficient in practice in an NVLINK-connected node, and hybrid SP-Ulysses and SP-Ring communication patterns sometimes bring better performance.
@@
We benchmark llama3-8B on two RDMA-connected 8xA800 nodes. Since LLAMA3-8B has only 8 heads, the maximum product of ulysses-degree and tp-degree is 8. Table 6 presents throughput in FLOPS on 64K, 80K and 120K.
@@
The optimal performance for sequence lengths of 64K and 80K is achieved with SP-only without TP-sp, the optimal setting of both is ulyess-degree=4 and a ring-degree=4. For sequence length 64K and 80K , the unified SP outperforms the SP-Ring by 13% and 12%, respectively. These results echo our conclusion in Tip 1.
For sequence length 64K and 80K, we can increase the global batch size (global-bs) of 2 using TP-sp by setting tp-degree=8, but it is not possible with SP-only. Increasing the global-bs to 2 can lead to 2.7% improvement in throughput to SP-only at the sequence length 64K. But at sequence length 80K, TP-sp+SP with global-bs as 2 is still worse than SP-only with global-bs as 2.
When the sequence length reaches 120K, the optimal performance is achieved with tp-degree=4 and ulysses-degree=2, reaching 266 TFLOPS, with an MFU of 0.85. At this sequence length, SP-only will encounter an Out-Of-Memory (OOM) issue. This confirms our Tip 5, which implies that the memory efficiency of SP is inferior to that of TP-sp. It is noteworthy that two other TP-SP hybrid settings yield similar performance, 0.82 and 0.84 in MFU, which is quite close to the optimal.
@@
We explored the upper bound of sequence length on 2 nodes, and the results are presented in Table 7. The largest sequence length was achieved with tp-degree=8 and ring-degree=2. At this point, SP-only could not run due to OOM. Compared to SP, TP-sp is more memory-efficient, hence for training the longest sequences, parallel degrees are limited by 8 attention head numbers should all be assigned to TP.
As the sequence length increases from 64K to 208K, the trend in MFU is a gradual improvement. This is due to the $O(L^{2})$ computational cost of attention quadratic increases, which results in a decreasing communication-to-computation ratio.
@@ 6 Future Work
As more and more models transition to a Mixture-of-Experts (MoE) architecture, the research on combining SP with expert parallelism (EP) becomes increasingly significant.
@@ 7 Conclusion
In this paper, we propose an approach that unifies DeepSpeed-Ulysses and Ring-Attention for sequence parallelism. This method encompasses the capabilities of both techniques, broadening the applicability and delivering superior performance in some cases. We systematically analyzed the interplay between sequence parallelism and other established parallelism methods, deriving a set of best practice conclusions. These conclusions have been validated through experimental data obtained from two GPU nodes, verifying the practical implications of our findings.

### v3

@@ Abstract
Sequence parallelism (SP), which divides the sequence dimension of input tensors across multiple computational devices, is becoming key to unlocking the long-context capabilities of generative AI models. This paper investigates the state-of-the-art SP approaches, i.e. DeepSpeed-Ulysses and Ring-Attention, and proposes a unified SP approach, which is more robust to transformer model architectures and network hardware topology. This paper compares the communication and memory cost of SP and existing parallelism, including data/tensor/zero/pipeline parallelism, and discusses the best practices for designing hybrid 4D parallelism involving SP. We achieved 86% MFU on two 8xA800 nodes using SP for sequence length 208K for the LLAMA3-8B model. Our code is publicly available on https://github.com/feifeibear/long-context-attention.
@@ 3 Unified Ulysses-Ring Sequence Parallelism
SP-Ulysses is sensitive to the number of attention heads. The parallelism degree  [^1] The parallelism degree is the number of devices that participate in parallel computation. In other words, it is the number of processes in a parallel process group. of DS-Ulysses cannot exceed the number of attention heads $hc$. Consequently, it is not suitable for GQA (Grouped Query Attention) [10] and MQA (Multi-Query Attention) [11] scenarios. For instance, Llama3-8B employs GQA with a KV head number of 8, which means that when using DS-Ulysses SP, the maximum SP degree is 8. However, if GQA is used and KV head number is 1, DS-Ulysses dose not function. In addition, since Tensor Parallelism also requires division across the $hc$ dimension, SP-Ulysses, and TP are in conflict.
@@ 4 SP in 4D Parallelism
As shown in Table 2, we analyze the communication and memory cost of a standard transformer block for different parallelism. The impact of GQA is not reflected in the table, but we will analyze it later. Entries in the Communications Params columns indicate the collective communication operations are conducted on the parameters and gradients of the transformer block. It includes the parameters/gradients of the weight and bias tensors of 4 Linear Layers in the Self-Attention layer, as well as 2 Linear Layers of the FFN layer, amounting to a volume of $12\times O(d^{2})$ elements in the GPT-2 model. Note that llama3 and llama2 has $12\times O(d^{2})$ elements since the intermediate size is lower. Entries in Communications Act columns indicate the collective communications are conducted on a single hidden states tensor belonging to activations, containing $bs\times L\times d$ elements. The memory cost is broken down into model parameters and gradients (P/G), Optimizer States (OS), and intermediate activation tensors (Act).
@@
Tip 3: We suggest that when utilizing SP, it should always be used in conjunction wit ZeRO-1/2. One can also consider employing ZeRO-3, and Offload techniques [5, 15] to trade off communication cost for memory savings.
@@
GQA/MQA can reduce SP communication costs, while the communication cost of TP-sp remains unchanged. Assuming the GQA group number is $G$, the Ulysses communication cost for $K$ and $V$ is reduced to 1/$G$, and the activation communication cost is reduced to $4O(bs*L*d)+4O(bs*L*d/G)$. Additionally, SP-Ring, which only communicates $K,V$, also leads to a 1/$G$ reduction in P2P bandwidth demand.
Tip 4: We suggest that SP may have an advantage over TP when employing GQA in terms of communication cost, as GQA can reduce the communication cost of SP without affecting TP.
@@ 5.1 Performance of Unified SP
As shown in Table 3, on an 8xA100-SXM4 NVLink node, the highest throughput for both 32K and 128K sequence lengths was achieved when the ulysses-degree is set to 8, that is the same as SP-Ulysses. Ulysses demonstrated a significant advantage over SP-Ring. The lb-ring indicates the load balance Ring-Attention. It consistently outperforms the basic ring attention implementation in terms of efficiency, with the advantage becoming more pronounced as the sequence length increases. This verified the argument in Sec. 3 of SP-Ring, that although communication overhead can be hidden through overlapping with computation, it results in a reduction in computational efficiency.
@@ 5.2 End-to-end SP Performance in Megatron-LM
Firstly, we compare the performance of SP and DP under the same LLAMA2-7B workload on a single node of A800 GPUs. The global batch size is 8. We use the SP-Unified, and pick the best ulysses-degree and ring-degree settings. In a single node, ulysses-degree as 8 usually achieves the best performance. As shown in Figure 5, DP outperforms SP-Unified across various input sequence lengths, which confirms our conclusion in Tip 2.
@@ 5.4 Hybrid SP and TP
Table 5 presents the performance of the llama2-7B model on a single node of 8xA800 GPU 80GB. The longest sequence length that can be achieved is 64K, and the best performance is obtained when tp-degree=4 and ulysses-degree=2. It outperforms TP-sp-only by 10%. When seqlen=64K, SP-only will bring OOM issue, which echos our conclusion in Tip 5 that SP is less memory-efficient than TP-sp.
When global-bs=16 and seqlen=30K, SP-Ulysses delivers the optimal performance, significantly surpassing other SP and TP hybrid strategies. It is also 26% better than TP-sp-only in throughput. This indicates that, despite the similar communication cost of SP-Ulysses and TP-sp, SP-Ulysses is more communication efficient than TP-sp in practice in an NVLINK-connected setting, and communication patterns of hybrid SP-Ulysses and SP-Ring sometimes bring better performance.
@@
We employed a hybrid parallel strategy combining TP and SP-Unified to benchmark the training throughput of LLAMA3-8B across two nodes. The results are shown in Table 6. In most of cases, the global batch size is fixed to be 1, aiming to maximize the sequence length. Since llama3-8B has only 8 heads, the maximum product of ulysses-degree and tp-degree is 8. Table 6 presents throughput in FLOPS on 64K, 80K and 120K.
@@
The optimal performance for sequence lengths of 64K and 80K is achieved with SP-only without TP-sp, the optimal setting of both is ulysses-degree=4 and ring-degree=4. For sequence length is 64K and 80K, the unified SP outperforms the SP-Ring by 13% and 12%, respectively. These results echo our conclusion in Tip 1.
For sequence length are 64K and 80K, we can increase the global batch size (global-bs) for TP+SP hybrid setting when tp-degree=8. However, SP-only always has OOM issue with global-bs=2. Increasing the global-bs to 2, TP+SP has 2.7% improvement in throughput to SP-only at the sequence length 64K. But at sequence length 80K, TP+SP with global-bs as 2 is still worse than SP-only with global-bs as 1.
When the sequence length reaches 120K, the optimal performance is achieved with tp-degree=4 and ulysses-degree=2, reaching 266 TFLOPS, with an MFU of 0.85. At this sequence length, SP-only meets an OOM issue. This confirms our Tip 5 again, which implies that the memory efficiency of SP is inferior to that of TP-sp. It is noteworthy that two other TP+SP hybrid settings yield similar performance, 0.82 and 0.84 in MFU, which is quite close to the optimal.
@@
We explored the upper bound of sequence length on 2 nodes, and the results are presented in Table 7. The largest sequence length was achieved with tp-degree=8 and ring-degree=2. At this point, SP-only could not run due to OOM. Compared to SP, TP-sp is more memory-efficient, hence for training the longest sequences, parallel degrees are limited by 8 attention head numbers should all be assigned to TP-sp.
As the sequence length increases from 64K to 208K, MFU is generally improving. This is due to the $O(L^{2})$ computational cost of attention quadratic increases, which results in a decreasing communication-to-computation ratio.
@@ 6 Future Work
As more and more models transition to a Mixture-of-Experts (MoE) architecture, the research on combining SP with expert parallelism (EP) becomes increasingly significant. Since TP has been proven to be better replaced by EP (Expert Parallelism) [18] due to its inferior communication cost, we argue that SP should also be given a lower priority than EP. This paper does not conduct experiments on MoE models, but we plan to test it in the future.
@@ 7 Conclusion
In this paper, we propose an approach that unifies DeepSpeed-Ulysses and Ring-Attention for sequence parallelism. This method encompasses the capabilities of both techniques, broadening the applicability and delivering superior performance in some cases. We systematically analyzed the interplay between sequence parallelism and other established parallelism methods, deriving a set of best practice suggestions. These suggestions have been validated through experimental results obtained from two GPU nodes.